# Deep Discriminative to Kernel Density Graph for In- and Out-of-distribution Calibrated Inference

### v7

@@ 5.2 Empirical Study on Vision Data
In vision data, each image pixel contains local information about the neighboring pixels. To extract the local information, we use the convolutional layers at the front-end of a convolutional neural network (CNN). More precisely, we have a convolutional front-end encoder, $h_{e}:\mathbb{R}^{D}\mapsto\mathbb{R}^{m}$ and typically, $m<<D$. After the encoder there is a few fully connected dense layers for discriminating among the $K$ class labels, $h_{f}:\mathbb{R}^{m}\mapsto\mathbb{R}^{K}$. We fit a KDN on the $m$-dimensional embedding outputs from the encoder and measure our proposed kernel in Equation (16) on the dense layers.
@@ 5.2.1 Simulation Study
Figure 3: KDN filters out inference points with different kinds of semantic shifts from the training data. Simulated images used as training points: (A) circle with radius $10$, (B) rectangle with sides $(20,50)$ and out-of-distribution test points: (C) ellipse with minor and major axis $(10,30)$. For the above images, RGB intensities are fixed at $[127,127,127]$. Mean max confidence of KDN are plotted for semantic shift of the inference points created by (D) changing the color intensity, (E) taking convex combination of circle and rectangle, i.e., $\epsilon\times\text{circle}+(1-\epsilon)\times\text{rectangle}$, (F) changing one of the axes of the ellipse. Note that for $b=10$, the ellipse becomes a circle (ID).
For the simulation study, we use a simple CNN with one convolutional layer ($3$ channels with $3\times 3$ kernel) followed by two fully connected layers with $10$ and $2$ nodes in each. We train the CNN on $2000$ circle (radius $10$) and $2000$ rectangle (sides $20,50$) images with their RGB values being fixed at $[127,127,127]$ and their centers randomly sampled within a square with sides $100$. The other pixels in the background where there is no object (circle, rectangle or ellipse) were set to 00.
We perform three experiments while inducing semantic shifts in the inference points as shown in Figure 3. In the first experiment, we randomly sampled data similar to the training points. However, we added the same shift to all the RGB values of an inference point (shown as color intensity in Figure 3 D). Therefore, the inference point is ID for color intensity at $127$ and otherwise OOD. In the second experiment, we kept the RGB values fixed at $[127,127,127]$ while taking convex combination of a circle and a rectangle. Let images of circles and rectangles be denoted by $X_{c}$ and $X_{r}$. We derive an interference point as $X_{inf}$:
$X_{inf}=\epsilon X_{c}+(1-\epsilon)X_{r}$ (20)
Therefore, $X_{inf}$ is maximally distant from the training points for $\epsilon=0.5$ and closest to the ID points at $\epsilon=0,1$. In the third experiment, we sampled ellipse images with the same RGB values as the training points. However, this time we gradually change one of the ellipse axes from $0.01$ to $40$ while keeping the other axis fixed at $10$. As a result, the inference point becomes ID for the axis length of $10$. As shown in Figure 3 (D, E, F), in all the experiments KDN becomes less confident for the OOD points while the parent CNN remains overconfident throughout the semantic shifts of the inference points. The above property shown by KDN is highly desirable in a learning algorithm as the model is resonant or tuned to the training inputs and suppresses anything that deviates from the training concepts.
@@ 5.2.2 Vision Benchmark Datasets Study
Table 1: KDN achieves better calibration compared to different algorithms on different datasets when trained on CIFAR-10. $\uparrow$’ and $\downarrow$’ indicate whether higher and lower values are better, respectively. Bolded indicates most performant, or within the margin of error of the most performant.
@@
Dataset Statistics Parent KDN Isotonic Sigmoid ACET ODIN OE

### v9

@@ 3.2 Empirical Study on Vision Data
In vision data, each image pixel contains local information about the neighboring pixels. To extract the local information, we use convolutional or vision transformer encoders at the front-end. Consider a set of images $\mathcal{I}$ where each image has width $W$, height $H$ and each pixel in an image is identified by $\{\mathbf{M},w_{1},w_{2}\}$, where $\boldsymbol{M}$ is a vector representing the RGB pixel values, $w_{1}$ is the horizontal location, $w_{2}$ is the vertical location. More precisely, we have a front-end encoder, $h_{e}:\mathcal{I}\mapsto\mathbb{R}^{d}$ and typically, $d<<W\times H$. After the encoder, there are a few fully connected dense layers for discriminating among the $K$ class labels, $h_{f}:\mathbb{R}^{d}\mapsto\mathbb{R}^{K}$. Note that the $d$-dimensional embedding outputs from the encoder are partitioned into polytopes by the dense layers (see Equation (5)) and we fit a KDN on the embedding outputs.
Table 1: KDN achieves good calibration at both ID and OOD regions whereas other approaches excel either in the ID or the OOD region. Notably, KDN has reduced confidence on wrongly classified ID points.$\uparrow$’ and $\downarrow$’ indicate whether higher and lower values are better, respectively. OCE = OOD calibration error, MMC = Mean Max Confidence on wrongly classified ID points.
(a) In-Distribution (ID) Performance
ID dataset Statistics Parent KDN Isotonic Sigmoid ACET ODIN OE Focal textbfDual Focal
[…]
ECE $\downarrow$ $0.01\pm 0.00$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.00}\pm 0.00$ $0.01\pm 0.00$ $0.02\pm 0.00$ $0.01\pm 0.00$ $0.07\pm 0.00$ $0.01\pm 0.00$
MMC $\downarrow$ $0.76\pm 0.01$ $\mathbf{0.65}\pm 0.08$ $0.74\pm 0.02$ $0.90\pm 0.01$ $0.86\pm 0.02$ $0.97\pm 0.01$ $0.69\pm 0.01$ $0.56\pm 0.01$ $0.69\pm 0.01$
[…]
ECE $\downarrow$ $0.10\pm 0.01$ $\mathbf{0.02}\pm 0.00$ $0.04\pm 0.01$ $0.02\pm 0.00$ $0.11\pm 0.01$ $0.04\pm 0.00$ $0.03\pm 0.00$ $0.07\pm 0.00$ $0.02\pm 0.00$
MMC $\downarrow$ $0.46\pm 0.02$ $\mathbf{0.41}\pm 0.02$ $0.65\pm 0.03$ $0.45\pm 0.02$ $0.84\pm 0.03$ $0.77\pm 0.01$ $0.52\pm 0.02$ $0.46\pm 0.00$ $0.55\pm 0.00$
[…]
ECE $\downarrow$ $0.01\pm 0.00$ $\mathbf{0.00}\pm 0.00$ $0.01\pm 0.00$ $0.06\pm 0.00$ $0.05\pm 0.01$ $0.06\pm 0.01$ $0.01\pm 0.00$ $0.06\pm 0.00$ $0.07\pm 0.01$
MMC $\downarrow$ $0.82\pm 0.01$ $0.63\pm 0.00$ $0.67\pm 0.01$ $0.78\pm 0.00$ $0.88\pm 0.01$ $0.97\pm 0.01$ $0.70\pm 0.04$ $0.33\pm 0.00$ $\mathbf{0.41}\pm 0.00$
(b) Out-of-Distribution (OOD) Performance (OCE $\downarrow$)
ID dataset OOD dataset Parent KDN Isotonic Sigmoid ACET ODIN OE Focal Dual Focal
CIFAR-10 CIFAR-100 $0.47\pm 0.01$ $\mathbf{0.12}\pm 0.01$ $0.47\pm 0.01$ $0.69\pm 0.01$ $0.57\pm 0.01$ $0.79\pm 0.00$ $0.29\pm 0.01$ $0.47\pm 0.00$ $0.51\pm 0.00$
SVHN $0.87\pm 0.00$ $\textbf{0.01}\pm 0.00$ $0.85\pm 0.00$ $0.69\pm 0.01$ $0.90\pm 0.00$ $0.87\pm 0.00$ $0.04\pm 0.01$ $0.40\pm 0.04$ $0.54\pm 0.03$
Noise $0.28\pm 0.08$ $0.03\pm 0.02$ $0.30\pm 0.04$ $0.56\pm 0.12$ $\mathbf{0.01}\pm 0.00$ $0.53\pm 0.09$ $0.07\pm 0.02$ $0.52\pm 0.11$ $0.79\pm 0.00$
CIFAR-100 CIFAR-10 $0.23\pm 0.01$ $\mathbf{0.12}\pm 0.01$ $0.50\pm 0.04$ $0.22\pm 0.01$ $0.62\pm 0.04$ $0.54\pm 0.01$ $0.28\pm 0.02$ $0.31\pm 0.00$ $0.42\pm 0.00$
SVHN $0.20\pm 0.01$ $\mathbf{0.12}\pm 0.01$ $0.47\pm 0.06$ $0.20\pm 0.02$ $0.58\pm 0.05$ $0.57\pm 0.02$ $0.20\pm 0.02$ $0.21\pm 0.03$ $0.30\pm 0.03$
Noise $0.05\pm 0.02$ $\mathbf{0.03}\pm 0.00$ $0.38\pm 0.04$ $0.07\pm 0.06$ $0.08\pm 0.01$ $0.18\pm 0.06$ $0.07\pm 0.04$ $0.04\pm 0.03$ $\mathbf{0.03}\pm 0.01$
SVHN CIFAR-10 $0.40\pm 0.04$ $\mathbf{0.14}\pm 0.01$ $0.37\pm 0.02$ $0.53\pm 0.03$ $0.67\pm 0.04$ $0.80\pm 0.02$ $\mathbf{0.14}\pm 0.01$ $0.31\pm 0.00$ $0.40\pm 0.00$
CIFAR-100 $0.40\pm 0.03$ $\mathbf{0.14}\pm 0.01$ $0.37\pm 0.03$ $0.54\pm 0.03$ $0.66\pm 0.03$ $0.80\pm 0.02$ $0.16\pm 0.01$ $0.40\pm 0.00$ $0.48\pm 0.00$
Noise $0.36\pm 0.08$ $0.12\pm 0.04$ $0.32\pm 0.06$ $0.48\pm 0.06$ $\mathbf{0.03}\pm 0.01$ $0.67\pm 0.05$ $0.15\pm 0.03$ $0.24\pm 0.03$ $0.32\pm 0.04$
@@ 3.2.1 Simulation Study
For the simulation study, we use a CNN with one convolutional layer (kernel size $3\times 3$) followed by two fully connected layers with $10$ and $2$ nodes in each. Each experiment is repeated $10$ times to get the error bars. Consider two random variables $C_{1},C_{2}\sim U(0,100)$. We sample $2000$ circles $X_{c}$ and $2000$ rectangles $X_{r}$ grayscale training images using Equation 22 and 23 respectively. Each image has width $100$ and height $100$. The CNN is trained to classify an image into a circle or a rectangle.
$M=\begin{cases}127,&~{}\text{if}~{}(w_{1}-C_{1})^{2}+(w_{2}-C_{2})^{2}\leq 100\\ 0,&~{}\text{otherwise}.\end{cases}$ (22)
$M=\begin{cases}127,&~{}\text{if}~{}C_{1}-10\leq w_{1}<C_{1}+10\\ &\text{and}~{}C_{2}-25\leq w_{2}<C_{2}+25\\ 0,&~{}\text{otherwise}.\end{cases}$ (23)
We perform three experiments while inducing different types of shifts in the inference points (Figure 5). In the first experiment, we sample inference points using Equation 22 and 23, but sweep over different intensities for the object (Figure 5 D). Therefore, the inference point is maximally ID for intensity at $127$ and becomes more OOD as we move away from $127$. In the second experiment, we kept the intensity at $127$ while taking convex combination of a circle and a rectangle. Let images of circles and rectangles be denoted by $X_{c}$ and $X_{r}$. We derive an interference point as $X_{inf}$:
$X_{inf}=\epsilon X_{c}+(1-\epsilon)X_{r}.$ (24)
Therefore, $X_{inf}$ is maximally distant from the training points for $\epsilon=0.5$ and closest to the ID points at $\epsilon=\{0,1\}$ (Figure 5 E).
In the third experiment, we sampled ellipse images using Equation 25, sweeping $a$ from $0.01$ to $40$ (Figure 5 F). The inference point becomes ID at $a=10$ and remains OOD otherwise. As shown in Figure 5 (D, E, F), in all the experiments KDN becomes less confident for the OOD points while the parent CNN remains overconfident throughout different shifts of the test points.
$M=\begin{cases}127,&~{}\text{if}~{}\frac{(w_{1}-C_{1})^{2}}{a^{2}}+\frac{(w_{2}-C_{2})^{2}}{100}\leq 1\\ 0,&~{}\text{otherwise}.\end{cases}$ (25)
@@ 3.2.2 Vision Benchmark Datasets Study
In this study, we use ViT_B16 (provided in keras-vit package) vision transformer encoders (Dosovitskiy et al., 2020) pretrained on ImageNet (Deng et al., 2009) dataset and fine-tuned on the ID datasets. As described in Section 3.2, we use a ViT_B16 encoder and use Equation 12 on the last two fully connected layers to calculate the geodesic kernel. We use the same encoder for all the baseline algorithms and fine-tune it without freezing any weight. We experiment with popular benchmark datasets– CIFAR10, CIFAR-100 and SVHN. For each experiment, we train on one of the three datasets and test KDN for OOD performance on the other two datasets. We also consider a simple OOD dataset, noise, where we sample noise OOD samples of size $32\times 32\times 3$ according to a Uniform distribution with intensities within range $[0,1]$. Although noise samples is easier to detect as OOD point, we observe that many of the baseline approaches perform poorly even on the noise samples. Each experiment is repeated for $5$ different seeds to get an error bar. Table 1 shows that pretrained vision transformers are already well-calibrated for ID, and the OOD approaches (ACET, ODIN, OE) degrade ID calibration of the parent model. On the contrary, ID calibration approaches (Isotonic, Sigmoid) perform poorly compared to that of KDN in the OOD region. KDN achieves a compromise between ID and OOD performance while having reduced confidence on wrongly classified ID samples (see MMC in Table 1). See Appendix Table 2 for traditionally used statistics for OOD detection.
4 Discussion
4.1 Related Works and Our Contributions
Recently, Liu et al. (2023) showed that a deep network aware of distance (distance from training points) is able to capture the uncertainty associated with its prediction better than traditional approaches. However, the concept of distance awareness has previously been used by a number of approaches in the literature to detect OOD points. These approaches attempt to learn a generative model in the projected latent space and control the likelihoods far away from the training data. For example, Ren et al. (2019) used the likelihood ratio test to detect OOD samples. Wan et al. (2018) modified the training loss so that the downstream projected features follow a Gaussian distribution. In contrast to the above approaches, Liu et al. (2023) proposed an in-training approach which learns a map from the input space to a latent space and trains a Gaussian process model on top of that. Distance awareness from training points improves OOD detection and the Gaussian process improves ID calibration. However, OOD detection works as long as there are two distinguishable score sets for ID and OOD points, whereas calibration aims at estimating the true predictive uncertainty of these points (see Appendix A for details). In our work, we propose a simple post-hoc calibration approach using geodesic distance and achieve the aforementioned distance awareness from the training data by replacing the affine function over the nearest polytope with a Gaussian kernel. The proposed approach works on both random forests and deep networks. We validated our proposed approach on both tabular and vision datatsets, whereas most of the existing methods are tailor-made for vision problems.
4.2 Conclusions
In this paper, we demonstrate a simple intuition that transforms traditional deep discriminative models into a type of binning and kerneling approach. The bin boundaries are determined by the internal structure learned by the parent approach and Geodesic distance encodes the low-dimensional structure learned by the model. Moreover, geodesic distance introduced in this paper may have broader impact on understanding the internal structure of the deep discriminative models, which may be of interest for future work. Our code, including the package and experiments in this manuscript, will be made publicly available upon acceptance of the paper. Appendix Section A answers a number of frequently asked questions about this work.
Acknowledgements
The authors thank the support of the NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (NSF grant 2031985) and THEORINET. This work is graciously supported by the Defense Advanced Research Projects Agency (DARPA) Lifelong Learning Machines program through contracts FA8650-18-2-7834 and HR0011-18-2-0025. Research was partially supported by funding from Microsoft Research and the Kavli Neuroscience Discovery Institute. We also thank the helpful discussion with Will LeVine, Tyler M. Tomita, Ali Geisa, Tiffany Chu and Jacob Desman.
Appendix A Frequently Asked Questions
We received questions while presenting the proposed approach in different venues, and as we wish to reserve the main text for the core contents, we have compiled the explanation to these questions here.
What is the relation between OOD detection and OOD confidence calibration?
Our work addresses both in- and out-of-distribution (ID and OOD) calibration. While traditional ID calibration methods like Isotonic and Sigmoid regression focus on achieving calibrated inference within the ID region, they do not address OOD calibration. Conversely, there is another group of literature which is primarily concerned about detecting OOD points. These approaches such as ACET, OE, and ODIN mainly focus on OOD detection rather than OOD calibration. Calibration is harder than detection, akin to how regression is harder than classification. To see that calibration is harder than detection, consider the fact that a well-calibrated model can perform detection, but a model capable of detecting OOD points may not be calibrated. OOD detection works as long as there are two distinguishable score sets for ID and OOD points, whereas calibration aims at estimating the true predictive uncertainty of these points. To our knowledge, only Meinke et al. (2021) explicitly addressed OOD calibration (Section 3 Theorem 1 in their paper), but they do not consider ID calibration. Our work treats calibration problems as a continuum between the ID and OOD regions rather than addressing them separately.
Isn’t this just another out-of-distribution detection approach?
One may argue that the experimental settings essentially reduce the task to a binary one: whether to output normal confidence for an ID sample or just a prior for an OOD sample. A binary OOD detector can also immediately output a confidence based on the prior if a sample is identified as OOD. However, without ID calibration, there is no reason to expect that confidence to be calibrated within the ID region. In the experimental setting, as one moves further away from the ID setting, an OOD calibrated classifier will output the prior. However, close to the ID setting, the OOD calibrated classifier will output a probability that the sample is in any given class. This is in contrast to an OOD detector, which, regardless of how close or far the data are from the ID data, if is OOD, it always effectively outputs the prior. We would say that our method subsumes OOD detection, because it also includes ID calibration, and OOD calibration. To our knowledge, there are no other papers demonstrating any algorithm with all these properties.
Isn’t this just another kernel density estimation (KDE) approach?
There are similarities between KDE and the proposed method. However, there is an indicator function in Equation 7 which is implemented using the geodesic distance proposed in the draft (which is absent in KDE), which makes KDN, KDF scale better with higher dimensions. Moreover, the center and bandwidth of the Gaussian kernels are estimated in a data-driven manner using the representation learned by the parent discriminative approach.
Using Gaussian kernel for posterior estimation is not novel. Isn’t this similar to Gaussian Process regression?
Our proposed approach utilizes the internal structure learned by the parent approach to build an estimate of the posterior and scale for really high dimensional datasets. However, it is folk’s knowledge that Gaussian Process does not scale beyond dimension $20$.
Number of polytopes grows exponentially with the size of network. Isn’t the approach computationally infeasible?
Many people were concerned about the runtime of our approach, possibly it could be an exponential function of the number of nodes. However, we did additional experiments (see Fig. 6) where we show training and testing time both are linear in the number of the nodes. Moreover, training time is only 200 seconds even when there are 40,000 nodes running on a MacBook Pro with an Apple M1 Max chip and 64 GB of RAM. The number of total polytopes in KDN is upper bounded by the training sample size as we only consider the polytopes populated by training data (see the first paragraph of Section 2.3 and Equation 7).
Figure 6: Scaling plots for KDN on datasets with different feature dimensions. The training sample size is fixed at $10$,$000$ and the number of nodes in each layer is increased in a four layer fully connected network. Total number of polytopes populated by the training data saturates to training size (left) while training (middle) as well as testing time (right) scales linearly with node number.
How does the proposed approach perform in term of runtime compared to the baseline approaches?
OOD calibration approaches such as ACET, OE and ODIN take about 2 days, an hour, 6 hours, respectively on GPUs. In-distribution calibration methods such as isotonic regression and sigmoid regression take a few minutes and use CPUs. Our approach addresses both ID and OOD calibration while taking a few minutes to train on CPUs, rather than GPUs. All the computations were performed for producing the results in Table 1 using a MacBook Pro with an Apple M1 Max chip and 64 GB of RAM.
Why is there no non-asymptotic performance guarantee?
Because of “No Free Lunch" theorem, it is impossible to provide a non-asymptotic guarantee that works for all training distributions. Hence, we provided extensive empirical experiments encompassing simulation, tabular and vision datasets to build intuition about the finite sample performance of the proposed approach.
Why did you use a new statistic called OCE for OOD calibration measure?
OCE measures OOD calibration according to Equation 4 and assumes that true class conditional priors for the datasets are known. On the contrary, ECE is used for measuring ID calibration and is a surrogate measure used when true posteriors are not known. For example, in Figure 3 where we know the true distribution, we have used Hellinger distance from the true posteriors instead of ECE. Note that measuring ECE requires accuracy which requires class labels and there is no class label available for the OOD samples in our setting.
Why is there no OOD baseline approach for the tabular benchmark datasets?
ACET, ODIN, OE are tailor-made for vision problems, therefore we can not run them on tabular data using the author provided codes. To the best of our knowledge, the tabular OOD method (Ren et al., 2019) that we found does OOD detection, not calibration. As it does not yield any posterior, we could not benchmark with the above method.
In Table 1, it seems OOD approaches including ACET and ODIN can’t even beat the parent model under OOD settings.
ACET and ODIN highly depend on the model architecture and the nature of the ID and OOD testsets. Moreover, they also depend on the OOD set used to train them. We used the OOD set used by the authors of the above algorithms. All these factors contribute to their inconsistent performance across datasets and model architecture. Tajwar et al. (2021) found a qualitatively similar result.
What would be the AUROC, FPR, i.e. metrics commonly used in OOD detection?
To save space in the main text, we have added AUROC and FPR in Appendix Table 2. Note that these scores are used for OOD detection. However, we are addressing both ID and OOD calibration (Equation 4 in the main text). OOD detection works as long as there are two distinguishable score sets for ID and OOD points, whereas calibration aims at estimating the true predictive uncertainty of these points, both in ID and OOD regions. That being said, a well-calibrated model can perform detection, but a model capable of detecting OOD points may not be calibrated.
Table 2: KDN has nearly similar or better AUROC and FPR to those of baseline OOD detection $\uparrow$’ and $\downarrow$’ indicate whether higher and lower values are better, respectively.
@@ Appendix A Frequently Asked Questions
Dataset Statistics Parent KDN Isotonic Sigmoid ACET ODIN OE Focal Dual Focal